OpenStudy (anonymous):

Use the appropriate sum or difference identity to find the exact value of cos (7pi/12)

OpenStudy (anonymous):

-.26 ?

OpenStudy (cwrw238):

cos (pi/3 + pi/4) = cos pi/3 cos pi/4 - sin pi/3 sinn pi/4 = 1/2 * 1 / sqrt2 - sqr3/2 * 1/sqrt2 = 1 / 2sqrt2 - sqrt3 / 2sqrt2 = (1 - sqrt3) / 2 sqrt2

OpenStudy (cwrw238):

- if my arithmetic is correct - ill check it by calculator

OpenStudy (anonymous):

(1 - sqrt3) / 2 sqrt2 isnt an option

OpenStudy (cwrw238):

yes - its correct

OpenStudy (anonymous):

hey if you click the attachment and look at #23, you will see that i dont have that option

OpenStudy (cwrw238):

ok

OpenStudy (cwrw238):

right - i think its there but written with a rational denominator multipl top and bottom by sqrt2: 1 - sqrt3 * sqrt2 = sqrt2 - sqrt6 ------- ---- ----------- 2 sqrt 2 * sqrt2 = 4 thats one of the options right?

OpenStudy (cwrw238):

its always tidier to rationalise the denom.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

are you a math teacher

OpenStudy (cwrw238):

i used to be - retired now - thats my cat and grandaughter on my pic

OpenStudy (anonymous):

oh ok. what part of UK do you live?

OpenStudy (cwrw238):

swansea

OpenStudy (anonymous):

kool. I have really appreciate all the help you given me

OpenStudy (cwrw238):

ur welcome r u from uk?

OpenStudy (anonymous):

nah I live in the states, but I'm somewhat familiar with England. I though thats where you were from

OpenStudy (cwrw238):

smart of you - by the way i write i suppose - i'm from wales - west part of uk - great place but it rains a lot. swansea is the home of dylan thomas - a poet who was popular in the states

OpenStudy (anonymous):

yeah I did quick research on swansea after you told me

OpenStudy (cwrw238):

well talk to you again - time for me to help my wife wash up!!

OpenStudy (anonymous):

ok take care